OpenStudy (timaashorty):

For positive real numbers, s,g, and t, the equation s= 1/2 gt^2 expresses s in terms of g and t. The variable t may be expressed as a function of s and g by the equation t = ?

OpenStudy (timaashorty):

answer choices:

OpenStudy (timaashorty):

\[A. \sqrt\frac{ 2s }{ g }\] \[B. \sqrt\frac{ g }{ 2s }\] \[C. \sqrt\frac{ s }{ 2g }\] \[D. \sqrt{s - \frac{ 1 }{ 2 }} g\] \[E.\sqrt{s-2g}\]

OpenStudy (anonymous):